# Post-detection inference for sequential changepoint localization

### v1

@@ A.3 Omitted proofs from Section 4
Appendix B Simulations for point estimates
@@
B.1 Known pre- and post-change distributions
In this experiment, $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$ For detecting the change, we use the usual cumulative sum (CUSUM) and Shiryaev-Roberts (SR) procedures Tartakovsky et al., (2014) as defined below:
$\tau_{\text{cusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak$ (B.1)
$\tau_{\text{sr}}=\inf\left\{\tau\in\mathbb{N}:\sum_{j=1}^{\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (B.2)
where $A$ is some fixed threshold. For our experiments, we set the value of $A$ at $10000$. Once the change is detected, we obtain $\hat{T}$, as defined in (2.7), based on observations up to the detection time. Fig. B.1 shows histograms of the values taken by $\hat{T}$ over $10000$ independent experiments.
Figure B.1: Histogram of $\hat{T}$, defined in (2.7), for known pre- and post-change over $10000$ independent experiments with $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$
B.2 Known pre-change and composite post-change distributions
@@
Here, we use the weighted CUSUM/SR detectors defined in (3.16) and (3.17). We considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. Fig. B.2 shows histograms of the values taken by $\hat{T}$ defined in (3.5) over $10000$ independent experiments.
Figure B.2: Histogram of $\hat{T}$, defined in (3.5), for known pre-change and unknown post-change over $10000$ independent experiments with $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta\neq 0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
B.3 Composite pre- and post-change distributions
@@
In this experiment, we considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=0.6,d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ Fig. B.3 shows histograms of the values taken by $\hat{T}$ defined in (4.6) over $10000$ independent experiments. It is crucial to note that the RIPr-based estimator has a higher bias but lower variance.
Figure B.3: Histogram of $\hat{T}$, defined in (4.6), for unknown pre- and post-change over $10000$ independent experiments with $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively.
Appendix C A heuristic method for composite pre- or post-change

### v2

@@ A.3 Omitted proofs from Section 4
Appendix B Another concrete example with Poisson distribution
B.1 Known pre-change and composite post-change
Here is $F_{0}$ is $\text{Pois}(\theta_{0})$ and the post-change model is $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\in\Theta_{1}\}$.
In this setting, we have the following example of a $1-\beta$ confidence sequence [2] for $\theta_{1}$
$\mathcal{S}^{\prime}_{n}=\{\theta:n\theta-\sum_{i=t}^{t+n-1}X_{i}\log\theta\leq\log\left(\frac{1}{\beta}\right)+\log I(c,c\theta)-\log I(n+c,\sum_{i=t}^{t+n-1}X_{i}+c\theta)\}\cap\Theta_{1},$ (B.1)
[…]
We fix $t\in\{1,\cdots,\tau\}$. Let $\lambda_{t}=\max\{\theta_{0},\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$. We only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Pois}(\lambda_{t})$ and $U_{n,1}^{j},\cdots,U_{n,X_{n}^{j}}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Uniform}(0,1)$, independently for each $n\in\mathbb{N},j=1,\cdots,B$. For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta_{0}/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ otherwise. }\end{cases}$ (B.2)
Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta_{0})$, for $n<t$ and $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta)$, for $\theta\in\mathcal{S}^{\prime}_{\tau-t+1},n\geq t$.
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (2.8), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let $\tau^{\theta}_{j}$ be the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark B.1) and they are finite. So, we can choose $L=\infty$.
If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,V_{t,\infty}^{j}$ can be simplified as below:
$V_{t,\infty}^{j}$ $=\max_{\begin{subarray}{c}i,t^{\prime}\in\mathbb{N}:1\leq i\leq t^{\prime},\\ \max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}\end{subarray}}\max_{\theta\in{S}^{j}_{t,t^{\prime}}}\frac{L_{i}\left(\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-i+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}{L_{t}\left(\frac{\sum_{k=t}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-t+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}.$
Here ${S}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},t\leq n\leq t^{\prime},\inf\mathcal{S}^{\prime}_{\tau-t+1}\leq\lambda_{t}U_{n,i}^{j}\leq\sup\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\inf\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$ is the set consisting of one candidate element from is a finite set with at most $2+\sum_{n=t}^{t^{\prime}}X_{n}^{j}$ many elements. All the maximum are taken over a finite set; hence, $V_{t}^{j}$ can be computed easily.
Remark B.1.
For weighted CUSUM or SR detectors (defined in (3.15),(3.16)), finding $t_{1}^{j},t_{2}^{j}$ is often straightforward. For example, consider $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{(\tau-j+1)\times(\theta-\theta_{0})\}(\theta/\theta_{0})^{\sum_{i=j}^{\tau}X_{i}}$ increases as $X_{i}$’s increases and hence both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\neq\theta_{0}\}$, if we employ Gamma weights (it is a conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
B.2 Composite pre- and post-change
Here we assume the pre and post-change model to be $\mathcal{P}_{i}=\{\text{Pois}(\theta):\theta\in\Theta_{i}\}$, for $i=0,1$ respectively.
In this setting, we have the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (B.1) and the same method can be employed for constructing $1-\gamma$ confidence interval for $\theta_{0}$ as well.
We fix $t\in\{1,\cdots,\tau\}$. Let $\lambda_{t}=\max\{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$. We only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Pois}(\lambda_{t})$ and $U_{n,1}^{j},\cdots,U_{n,X_{n}^{j}}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Uniform}(0,1)$, independently for each $n\in\mathbb{N},j=1,\cdots,B$. For $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$,
$Y_{n}^{j}(\theta,\theta^{\prime}):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta^{\prime}/\lambda_{t}),\text{ otherwise. }\end{cases}$ (B.3)
Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta)$, for $n<t$, $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta^{\prime})$, for $n\geq t$ and $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (4.8), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. For Poisson distribution as well, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly for the same detection algorithms, similar to the Gaussian setting discussed in Remark 4.5.
If $t_{2}^{j}<t$, we have $U_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified as below:
$U_{t,\infty}^{j}$ $=\max_{\begin{subarray}{c}i,t^{\prime}\in\mathbb{N}:1\leq i\leq t^{\prime},\\ \max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}\end{subarray}}\max_{\begin{subarray}{c}\theta\in{A}^{i,j}_{t,t^{\prime}}\\ \theta^{\prime}\in{B}^{i,j}_{t,t^{\prime}}\end{subarray}}\frac{L_{i}\left(\frac{\sum_{k=1}^{i-1}Y_{k}^{j}(\theta,\theta^{\prime})}{i-1},\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime})}{t^{\prime}-i+1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)}{L_{t}\left(\frac{\sum_{k=1}^{t-1}Y_{k}^{j}(\theta,\theta^{\prime})}{t-1},\frac{\sum_{k=t}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime})}{t^{\prime}-t+1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)}.$
Here ${A}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},n\leq t-1,\inf\mathcal{S}_{t-1}\leq\lambda_{t}U_{n,k}^{j}\leq\sup\mathcal{S}_{t-1}\}\cup\{\inf\mathcal{S}_{t-1}\}\cup\{\sup\mathcal{S}_{t-1}\}$ and ${B}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},t\leq n\leq t^{\prime},\inf\mathcal{S}^{\prime}_{\tau-t+1}\leq\lambda_{t}U_{n,k}^{j}\leq\sup\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\inf\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$ are finite sets with at most $(2+\sum_{n=1}^{t^{\prime}}X_{n}^{j})$ many elements. All the maximum are taken over a finite set; hence, $U_{t}^{j}$ can be computed easily.
B.3 Experimental results
B.3.1 Known pre- and post-change (Setting I)
We consider Poisson the mean-change problem: $\text{Pois}(1)$ (pre-change) vs $\text{Pois}(2)$ (post-change) with the true (unknown) changepoint at $T=100,500$. We employ the CUSUM change detection method defined in (C.1) with thresholds at $A=1000,10000$. After detection, we employ Algorithm 2 to construct confidence sets for $T$. 5.1 and B.1 provide visualizations of the confidence sets and point estimates (2.7) of the changepoints across $5$ runs.
Figure B.1: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois$(1)$ and the rest from Pois$(2)$, with $T=100$ (top row) or $T=500$ (bottom row). Change is detected using CUSUM detector (C.1) with $A=10000$ (top row) or $A=1000$ (bottom row). The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points. $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table B.2.
Table B.2: Setting I: Average (of $100$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (2.7) and average detection delay (given a true detection) of the detection algorithm used (C.1) with threshold $A$, while varying $T$ and $A$.
$T$ $A$
Conditional
coverage
Unconditional
coverage
100 10000 0.94 0.94 18.04 4.02 22.15
100 1000 0.94 0.94 17.45 3.98 15.72
500 10000 0.95 0.95 18.53 4.30 22.10
500 1000 0.96 0.89 18.72 3.96 16.14
B.3.2 Known pre-change and composite post-change (Setting II)
We perform experiments for Poisson rate-change scenarios having pre-change distribution as $\text{Pois}(1)$, and the true (unknown) post-change data distribution as $\text{Pois}(2)$, with the changepoint at $T=100,200,$ assuming the composite post-change model to be $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$. For detection, we employ a weighted CUSUM detector (3.15) with thresholds at $A=1000,10000$. For weighted CUSUM, we employ a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, and $\theta=1.9$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we construct confidence sets (as defined in (3.11)) for $T$.
We report the average size and the conditional and unconditional coverage across $100$ independent runs in (5.4). The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table B.2). We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table. We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
Table B.3: Setting II: Average (of $100$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets (3.11) with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ (3.5) and average detection delay (given no false detection) of the change detection method used (3.15) with threshold $A$, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
T A
Conditional
coverage
Unconditional
coverage
100 10000 0.97 0.97 33.13 6.11 27.68
100 1000 0.97 0.96 33.85 5.59 19.14
200 10000 0.98 0.98 36.08 5.98 26.81
200 1000 0.97 0.93 35.24 5.74 21.27
B.3.3 Composite pre- and post-change (Setting III)
Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $200$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively. For change detection, we employ a weighted CUSUM-type detector (defined in (4.17)), where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, and for the post-change parameter, the same discrete weight distribution as in Section B.3.2 is considered.
After detection, we construct confidence sets (as defined in (4.12)) for $T$ (assuming $T\neq 1$). We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$ (defined in (2.7)), and delay (given no false detection) across $100$ independent runs in Table B.4.
Table B.4: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets (4.12) with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ (4.6) and average detection delay (given no false detection) of the change detection method used (4.17) with threshold $A$, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
T A
Conditional
coverage
Unconditional
coverage
100 10000 0.97 0.97 35.04 6.34 31.56
100 1000 0.97 0.97 35.19 6.27 26.33
200 10000 0.98 0.98 37.54 5.92 30.45
200 1000 0.98 0.97 37.86 6.18 25.82
However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table B.2). We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table B.2), exhibiting both higher coverage and larger size.
Appendix C Simulations for point estimates
@@
C.1 Known pre- and post-change distributions
In this experiment, $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$ For detecting the change, we use the usual cumulative sum (CUSUM) and Shiryaev-Roberts (SR) procedures [23] as defined below:
$\tau_{\text{cusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak$ (C.1)
$\tau_{\text{sr}}=\inf\left\{\tau\in\mathbb{N}:\sum_{j=1}^{\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (C.2)
where $A$ is some fixed threshold. For our experiments, we set the value of $A$ at $10000$. Once the change is detected, we obtain $\hat{T}$, as defined in (2.7), based on observations up to the detection time. Fig. C.1 shows histograms of the values taken by $\hat{T}$ over $10000$ independent experiments.
Figure C.1: Histogram of $\hat{T}$, defined in (2.7), for known pre- and post-change over $10000$ independent experiments with $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$
C.2 Known pre-change and composite post-change distributions
@@
Here, we use the weighted CUSUM/SR detectors defined in (3.15) and (3.16). We considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. Fig. C.2 shows histograms of the values taken by $\hat{T}$ defined in (3.5) over $10000$ independent experiments.
Figure C.2: Histogram of $\hat{T}$, defined in (3.5), for known pre-change and unknown post-change over $10000$ independent experiments with $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta\neq 0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
C.3 Composite pre- and post-change distributions
@@
In this experiment, we considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=0.6,d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ Fig. C.3 shows histograms of the values taken by $\hat{T}$ defined in (4.6) over $10000$ independent experiments. It is crucial to note that the RIPr-based estimator has a higher bias but lower variance.
Figure C.3: Histogram of $\hat{T}$, defined in (4.6), for unknown pre- and post-change over $10000$ independent experiments with $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively.
Appendix D A heuristic method for composite pre- or post-change